A car travels in straight line with an average velocity of 80 km/h for 2.5 h and then an average velocity of 40 km/h from 1.5 h (a) what is the total displacement for the 4 h trip? (b) What is the average velocity for the total trip?
1
Expert's answer
2019-03-19T12:07:27-0400

Question (a).

Total displacement:

"s = s_1 + s_2 = v_1t_1 + v_2t_2 = 80~\frac{\text{km}}{\text{h}}~*~2.5~\text{h} + 40~\frac{\text{km}}{\text{h}}~*~1.5~\text{h} = 260~\text{km}."

Question (b).

Average velocity for the total trip:

"v_\text{avg} = \frac{s}{t} = \frac{v_1t_1 + v_2t_2}{t_1+t_2} = \frac{260~\text{km}}{2.5~\text{h} + 1.5~\text{h}} = 65~\frac{\text{km}}{\text{h}}."